OpenStudy (anonymous):

Robbie the Robot is on a weather satellite orbiting Earth about 3600 km above the surface. The Earth’s radius is about 6400 km. He has had a malfunction in his output device, and the satellite is traveling without communication. His last report was only in terms of trigonometric values and was only partially received. It said, sin Θ < 0….. and then he was lost again.

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

Okay, whats the question ?

OpenStudy (anonymous):

Part 1: Create a set of coordinates that would be reasonable for Robbie’s position in space and satisfy his last, partial report. Using complete sentences, describe Robbie’s location and your reasoning. Part 2: What are the values of the sin Θ, cos Θ, and tan Θ using your coordinate point? @ganeshie8

ganeshie8 (ganeshie8):

Since Robbie said \(\sin \theta \lt 0\), he must be in III or IV quadrants where sin value is negative. Also since earth radius is 6400 km and he is 3600 km above the surface, that means he is 10000 km from the center of earth \(r = 10000\) \(\theta = -45^{\circ}\) as \(\theta = -45^{\circ}\) satisfies the Robie's given specification the coordinates will be : \((10000 \cos (-45), ~10000\sin(-45)) = \left(\frac{10000}{\sqrt{2}}, ~\frac{-1000}{\sqrt{2}}\right)\)

ganeshie8 (ganeshie8):

thats for part A ^

ganeshie8 (ganeshie8):

I'll let u work the part 2